Joe works at a local department store. His manager asked him to choose for display 5 different shirts

from the wall of the store that has 8 shirts on it to put in a display. How many ways can Joe choose the

shirts?


1
Expert's answer
2021-05-07T14:17:56-0400

Total number of shirts=8


Number of shirts joe has to display=5


So Number of ways joe can choose shirt-


"= ^8C_5=\dfrac{8!}{5!(8-5)!}=\dfrac{8\times 7\times 6\times 5!}{6\times 5!}=56\text{ ways}"